OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle whose legs have lengths of 5 and 12? 17?

OpenStudy (anonymous):

use a^2+b^2=c^2

OpenStudy (anonymous):

\[c^2 = a^2 + b^2\] c is the hypotenuse here..

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Let him try his best @mandonut

OpenStudy (anonymous):

are you in a quiz or test @helpme9283

OpenStudy (anonymous):

yeah you shouldnt give me the answer first but you should let me try it.. but thanks

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i'm working on practice problems

OpenStudy (anonymous):

Then it is okay..